Find the dimensions of a rectangle whose area s 575square cm and whose perimeter is 96cm
Answer by LinnW(1048)

set l = length
w = width
l*w = area
l*w = 575
2l + 2w = 96
solve for l
add -2w to each side
2l = 96 -2w
divide by 2
l = 48 -w
substitute 48 -w for l in l*w = 575
(48-w)(w) = 575
48w - w^2 = 575
add -575 to each side
-w^2 +48w -575 = 0
multiply by -1
w^2 -48w + 575 = 0
(w - 23)(w - 25) = 0
w = 23 or w = 25
Our two dimensions are 23 and 25
23*25 gives area
2(23) + 2(25) = 96